Write an equation in slope-intercept form of the line satisfying the given condition
the line passes through (4,-5) and is parallel to the line whose equation is y=3x+5
type the slope-intercept form of the equation of the line

Answer by mananth(16946)

1 y = 3 x + 5
Divide by 1
y = 3 x + 5
Compare this equation with y=mx+b
slope m = 3
The slope of a line parallel to the above line will be the same
The slope of the required line will be 3
m= 3 ,point ( 4 , -5 )
Find b by plugging the values of m & the point in
y=mx+b
-5 = 12 + b
b= -17
m= 3
Plug value of the slope and b in y = mx +b
The required equation is y=3x-17